# Early-Exit Graph Neural Networks

### v1

@@ 3 Methodology
Having addressed O1, we now turn to O2: implementing a contextual early-exit mechanism. We first focus on node classification; the extension to graph classification is presented in Appendix D. Recently, a very similar problem was faced by Finkelshtein et al. (2024), where each node had to take actions based on the context. Their strategy employed the straight-through Gumbel-Softmax estimator (Jang et al., 2017; Maddison et al., 2017), which derives a continuous approximation of discrete action sampling. Taking inspiration from Finkelshtein et al. (2024), we propose something similar but with a reduced action space. Let $\Omega$ be the action space, we have $|\Omega|=2$, where $\Omega=\{\texttt{0},\texttt{1}\}$, with 0 is the non-exiting, and 1 is the exiting action. In our work, we define $\mathbf{C}\in\mathbb{R}^{|\Omega|}$ as a confidence vector representing action probabilities. For example, $\mathbf{C}(\texttt{1})$ gives the probability of exiting at a specific network state. The Gumbel-Softmax estimator approximates the categorical distribution $\mathbf{C}$ using a Gumbel-distributed vector $\mathbf{g}\in\mathbb{R}^{|\Omega|}$, where each component $g(a)\sim\text{GUMBEL}(0,1)$ for $a\in\{\texttt{0},\texttt{1}\}$. Given $\mathbf{C}$ and a temperature parameter $\nu$, the Gumbel-Softmax score for node $i$ is computed as:
$\mathbf{c}_{i}(\mathbf{C}_{i};\nu_{i})=\nicefrac{{\exp\left(\frac{\log(\mathbf{C}_{i})+\mathbf{g}_{i}}{\nu_{i}}\right)}}{{\sum_{a\in\{\texttt{0},\texttt{1}\}}\exp\left(\frac{\log(C_{i}(a))+g_{i}(a)}{\nu_{i}}\right)}},$
we observe that as $\nu_{i}\to 0$, this expression approaches a one-hot encoding. We compute $\mathbf{C}_{i}$ and $\nu_{i}$ from the current context $\mathbf{H}^{t}$ in matrix form as $\mathbf{C}^{t}\in\mathbb{R}^{n\times|\Omega|}$ and $\bm{\nu}^{t}\in\mathbb{R}^{n\times 1}$, via $f_{c}$ and $f_{\nu}$ which are neural networks with fixed depth $L_{f}$ and hidden dimension $m_{f}$. For efficiency purposes, $f_{c}$ and $f_{\nu}$ are shared across layers, even though their outputs $\mathbf{C}^{t}$ and $\bm{\nu}^{t}$ depend on time $t$. Additional details on the Gumbel-Softmax distribution and implementation of $f_{c}$ and $f_{\nu}$ are provided in Appendix B.1. To the best of our knowledge, we are the first to implement early exit using the Gumbel-Softmax reparametrization trick. Following Finkelshtein et al. (2024), for node classification, we design $f_{c}$ and $f_{\nu}$ as GNNs to estimate $\mathbf{C}^{t}$ and $\bm{\nu}^{t}$. This choice allows each node to incorporate information from its $L_{f}$-hop neighborhood—effectively reasoning $L_{f}$-steps ahead when deciding whether to exit or adjust the temperature. We observe that the decision-making process is based on the context because of the hidden features, but also depending on the task. Indeed, the update of the weights of $f_{c}$ and $f_{\nu}$ is exclusively based on the gradients of the task loss (e.g., Cross-Entropy, Mean Squared Error), since we do not use additional losses.
Early-Exit Graph Neural Networks. As shown, SAS-GNN satisfies Goal O1. By combining it with the early-exit mechanism for Goal O2, we realize our complete framework called Early-Exit Graph Neural Networks (EEGNNs), and presented in Algorithm 1. The algorithm is designed for node classification on a single graph, but can be naturally extended to inductive settings by applying it per sample. The graph-level version is provided in Appendix D. To integrate our Early-Exit mechanism into SAS-GNN, we need to include the Gumbel-Softmax scores in the message-passing update. Finkelshtein et al. (2024) implements this by modifying the graph topology, meaning that the gradients of the task loss are computed w.r.t. $\mathbf{\bar{A}}$. In our case, we decided to use the integration constant $\tau$, which is no longer fixed, but it varies across the layers and also node-wise. This approach is novel in the literature, and we refer to it as Neural Adaptive-step Early-Exit. The idea is not only to make everything differentiable but, in general, by choosing $\bm{\tau}^{l}=\mathbf{c}^{l}(\texttt{0})$ as the probability of non-exiting, we keep updating the next node representation proportionally to this probability. In fact, if $\mathbf{c}^{l}(\texttt{0})=0$, then $\mathbf{H}^{l+1}\leftarrow\mathbf{H}^{l}$, which enforces further the early-exit bias. At each iteration, we save those nodes that are predicted to exit. At the end of the main loop, we use the last representations $\mathbf{h}_{i}^{L}$ to complete the output. In this algorithm, we have a different integration constant per node; for this reason, we view $l$ as the $l$-th exit point. $L$ can be seen as the number of candidate exit points. In the time domain, the total time that a node $u$ spends in the GNN is computed as $\sum_{l=0}^{L}\tau^{l}_{u}$, where each $\tau^{l}_{u}$ is predicted by a GNN and also corresponds with the non-exiting probability. In this sense, we can describe the exit of each node as continuous rather than limited to the discrete domain. While this remains beyond the scope of the paper and we do not investigate this aspect extensively, in Appendix F.11, we depict how each node can exit at a unique instant of time.
@@
To summarize, using SAS-GNN in line 6 of the algorithm offers two key benefits: (1) it provides a fallback that mitigates message-passing failures at any depth; and (2) weight sharing avoids parameter waste from unused layers.
Relation to Adaptive-Step in Runge-Kutta Solvers.Our Neural Adaptive-Step shares conceptual similarities with ODE solvers like DOPRI5/8 (Dormand and Prince, 1980), which also relax fixed-depth constraints. However, our approach differs in that a neural network determines the stopping condition, whereas DOPRI solvers rely on integration error tolerance. Although integrating these methods may be beneficial, it is outside the scope of this work.
Mitigating Under-reaching. EEGNN is inherently robust to under-reaching. We observe that setting a large $L$ does not hurt performance, as the model learns to exit early when appropriate. While related work learns depth during training (Errica et al., 2023), integrating this into our setting is left for future exploration.
@@
Discussion on Complexity. We analyze the space complexity in terms of parameter count, comparing MPNNs, GTs, and Co-GNNs with our models. Thanks to weight sharing, SAS-GNN maintains constant complexity w.r.t. $L$ and quadratic complexity in the hidden dimension $m^{\prime}$, as both $\mathbf{\Omega}_{as}$ and $\mathbf{W}_{s}$ are $m^{\prime 2}$ matrices. Their symmetry and antisymmetry reduce the number of unique parameters to approximately $\frac{1}{2}m^{\prime 2}\cdot 2$. EEGNN adds parameters via $f_{c}$ and $f_{\nu}$, but this overhead depends only on $L_{f}$, not $L$, since these modules are shared across layers. When implemented as SAS-GNNs, the overall complexity becomes $\mathcal{O}(m^{\prime 2}+m_{f}^{2})$. Table 1 also includes Polynormer, a GT-based model that uses two modules: local attention (with $L_{l}$ layers) and global attention (with $L_{g}$ layers), each using 4 non-shared weight matrices of size $m^{\prime 2}$. Assuming $L_{l}=L_{g}=L$, the total complexity becomes $\mathcal{O}(8Lm^{\prime 2})$ as shown in table. Regarding time complexity, SAS-GNN offers no clear runtime advantage, but EEGNN may reduce computation by exiting early for many nodes. Although some nodes may still require full-depth processing, potentially becoming bottlenecks, this is not always the case, as explored in Section 4.1.
@@ 4 Experimental Evaluation
In this section, we empirically validate the contributions of our proposed framework, with a particular emphasis on the multiple advantages offered by EEGNN, in terms of efficiency, effectiveness, modularity, and flexibility.
@@ 4.1 Impact of the Early-Exit Components
According to Algorithm 1, EEGNN, and its exit module (i.e., $f_{c}$, $f_{\nu}$) updates its weights directly through the task loss, conversely from previous methods, which typically require additional loss terms or regularizers to guide the learning process (Spinelli et al., 2021). This feature enables EEGNN to learn to exit in a task-driven fashion, providing a significant advantage over fixed-depth models, which otherwise require manual tuning of the optimal number of layers $L$ for each dataset and task.
To assess this, we evaluate on Peptides-func (graph classification) and Peptides-struct (graph regression), which share graph structures but differ in predictive goals (see Table 7 in the appendix), potentially requiring different problem radii. As shown in Figure 1, EEGNN effectively learns distinct exit distributions for each task. Notably, for Peptides-func, EEGNN predicts to exit after just two layers, suggesting long-range interactions are not essential for this task.
Our approach is also modular, owing to the use of the Straight-Through Gumbel-Softmax Estimator. By simply replacing line 6 in Algorithm 1, the early-exit mechanism can be integrated into other message-passing architectures. Here, parameter sharing is desirable for space efficiency, but it is not a hard constraint. To demonstrate this flexibility, in Appendix F.6, we attach the early-exit module to standard MPNNs and evaluate the combination across three datasets.
Results show that baseline models often degrade as depth increases, likely due to over-smoothing or over-squashing, while early exits help mitigate these issues, preserving or even improving performance, similarly to Spinelli et al. (2021). This highlights how depth, when not properly stabilized (e.g., through architectural choices like SAS-GNN), can hinder predictive accuracy. Later in this section, we show that equipping an already robust backbone like SAS-GNN with EEGNN does not aim to improve accuracy, but rather to enhance inference efficiency. We also find that replacing ReLU with a ReLU+TanH combination can help reduce degradation in some models.
Another substantial benefit of EEGNN is its ability to reduce inference time by skipping redundant computation for nodes that exit early. As shown in Table 2, we fix $m^{\prime}=32$ and vary $L$, averaging inference time over 1,500 forward passes. EEGNN and SAS-GNN maintain a constant number of parameters due to shared weights, unlike models such as GCN, Co-GNN, and Polynormer. While Co-GNN and Polynormer demonstrate strong performance on several benchmarks (Luo et al., 2024), they incur significantly higher computational and memory costs—Co-GNN due to expensive propagation mechanisms, and Polynormer due to its large parameter count. According to this analysis, EEGNN treats $L$ as a budget rather than a hyperparameter and effectively maintains nearly constant inference complexity. For instance, increasing $L$ does not involve a significant change to the inference time. For a more detailed runtime analysis, refer to Appendix F.9.
@@
Figure 1: EEGNN’s Graph exit layer distributions. Numbers on top of the bars correspond to the number of test graphs that were processed up to that layer. The associated performance is in Table 4.
Table 2: Runtime and parameter analysis in Amazon Ratings. Times are in seconds.
@@
Model Inference Time (s) Number of Parameters 10 Layers 20 Layers 10 Layers 20 Layers GCN $0.0269\pm 0.0113$ $0.0370\pm 0.0140$ 20,352 30,912 Co-GNN $0.0562\pm 0.0163$ $0.0838\pm 0.0262$ 35,049 55,849 Polynormer $0.0191\pm 0.0037$ $0.0310\pm 0.0046$ $47,306$ $80,266$ SAS-GNN $0.0278\pm 0.0106$ $0.0442\pm 0.0150$ 11,904 11,904 EEGNN $0.0267\pm 0.0128$ $0.0227\pm 0.0092$ 14,146 14,146
@@
Table 3: Node classification under heterophily. The scores are marked in red for the first, blue for the second, and green for the third. Scores are from Finkelshtein et al. (2024); Luo et al. (2024).

### v2

@@ 3 Methodology
Gumbel Softmax Early-Exit Mechanism (O2). Having addressed O1, we now turn to O2: implementing a contextual early-exit mechanism. We first focus on node classification; the extension to graph classification is presented in Appendix D. Taking inspiration from Finkelshtein et al. (2024), we employ the straight-through Gumbel-Softmax estimator (Jang et al., 2017; Maddison et al., 2017), which provides a differentiable relaxation of discrete sampling. Let $\Omega$ be the action space, $\Omega=\{\texttt{0},\texttt{1}\}$ that corresponds to continue (0) or exit (1). For each node, a confidence vector $\mathbf{C}\in\mathbb{R}^{|\Omega|}$ defines action probabilities (e.g., $\mathbf{C}(\texttt{1})$ is the exit probability). The Gumbel-Softmax estimator approximates the categorical distribution $\mathbf{C}$ using a Gumbel-distributed vector $\mathbf{g}\in\mathbb{R}^{|\Omega|}$, where each component $g(a)\sim\text{GUMBEL}(0,1)$ for $a\in\{\texttt{0},\texttt{1}\}$. For a node $i$, given its confidence vector $\mathbf{C}_{i}$ and a temperature parameter $\nu_{i}$, the Gumbel-Softmax score is computed as:
$\mathbf{c}_{i}(\mathbf{C}_{i};\nu_{i})=\text{Softmax}\left(\frac{\log(\mathbf{C}_{i})+\mathbf{g}_{i}}{\nu_{i}}\right)$
which approaches a one-hot encoding as $\nu_{i}\to 0$.
We compute $\mathbf{C}^{t}\in\mathbb{R}^{n\times|\Omega|}$ and $\bm{\nu}^{t}\in\mathbb{R}^{n\times 1}$ from hidden states $\mathbf{H}^{t}$ using two GNN modules, $f_{c}$ and $f_{\nu}$, with fixed depth $L_{f}$ and hidden size $m_{f}$. These are shared across layers for efficiency. To our knowledge, this is the first application of the Gumbel-Softmax reparametrization to early exit in GNNs. Importantly, $f_{c}$ and $f_{\nu}$ are trained end-to-end with only the task loss (e.g., cross-entropy, MSE), so exit decisions depend purely on task-driven context without auxiliary supervision. Additional details on the Gumbel-Softmax distribution and implementation of $f_{c}$ and $f_{\nu}$ are provided in Appendix B.1.
Early-Exit Graph Neural Networks. As shown, SAS-GNN satisfies Goal O1; combining it with the Gumbel-Softmax early-exit mechanism (O2) yields our full framework, EEGNNs, presented in Algorithm 1. The algorithm is presented for node classification but extends naturally to inductive and graph-level tasks (see Appendix D). To integrate our Early-Exit mechanism into SAS-GNN, we need to include the Gumbel-Softmax scores in the message-passing update. Unlike Finkelshtein et al. (2024), who modify the topology (and backpropagate through $\mathbf{\bar{A}}$), we adapt the integration constant $\tau$, making it node- and layer-dependent. We introduce so the Neural Adaptive-step: we set $\bm{\tau}^{l}=\mathbf{c}^{l}(\texttt{0})$, the non-exit probability. When $\mathbf{c}^{l}(\texttt{0})\to 0$, the update reduces to $\mathbf{H}^{l+1}\leftarrow\mathbf{H}^{l}$, naturally encoding the exit bias. Nodes predicted to exit are stored at step $l$, and any remaining nodes are output at depth $L$.
Thus, each node has a personalized trajectory: its total “time in the network” is $\sum_{l=0}^{L}\tau^{l}_{u}$, giving a continuous view of exit rather than discrete steps (visualized in Appendix F.14). Practically, $L$ acts as the maximum number of exit points. Setting $L$ high removes the need for manual depth tuning and enables long-range reasoning and deep feature processing, as long as intermediate features remain informative; setting it within hardware limits guarantees inference never exceeds the cost of a non-exiting model. Weight sharing ensures memory remains constant regardless of $L$.
@@
To summarize, using SAS-GNN in line 1 of the algorithm offers two key benefits: (1) it provides a fallback that mitigates message-passing failures at any depth; and (2) the weight sharing is a byproduct of the Neural ODE design, which is an advantage for space complexity, and also avoids parameter waste from unused layers. In Section F.10, we discuss the connection of our Neural Adaptive Step, with Adaptive Step in Runge-Kutta solvers (Dormand and Prince, 1980a), and also the mitigation of underreaching.
@@
Discussion on Complexity. We analyze the space complexity in terms of parameter count, comparing MPNNs, GTs, and Co-GNNs with our models. Thanks to weight sharing, SAS-GNN maintains constant complexity w.r.t. $L$ and quadratic complexity in the hidden dimension $m^{\prime}$ with symmetry and antisymmetry halving the effective number of parameters. EEGNN adds parameters via $f_{c}$ and $f_{\nu}$, but this overhead depends only on $L_{f}$, not $L$, since these modules are shared across layers. In the long-range experiments, we implemented $f_{c}$ and $f_{\nu}$ as additional SAS-GNNs, so there the overall complexity becomes $\mathcal{O}(m^{\prime 2}+m_{f}^{2})$. Table 1 also includes Polynormer, a GT-based model that uses two modules: local attention (with $L_{l}$ layers) and global attention (with $L_{g}$ layers), each using 4 non-shared weight matrices of size $m^{\prime 2}$. Assuming $L_{l}=L_{g}=L$, the total complexity becomes $\mathcal{O}(8Lm^{\prime 2})$ as shown in table. Regarding time complexity, SAS-GNN offers no clear runtime advantage, but EEGNN may reduce computation by exiting early for many nodes. Although some nodes may still require full-depth processing, potentially becoming bottlenecks, this is not always the case, as explored in Section 4.1.
@@ 4.1 Impact of the Early-Exit Components
Task-driven learning. Unlike approaches that rely on auxiliary budget-aware losses (Spinelli et al., 2021), EEGNN updates the exit network’s weights coming from ($f_{c},f_{\nu}$), directly via the task loss. This theoretically ensures that decisions are strictly context-aware and eliminates the need to tune depth $L$ per dataset. We emprically validate the efficacy of this task-driven mechanism through two sets of experiments.
Task awareness. We first evaluate the ability to recognize when deep computation is strictly necessary using the SSSP dataset from the ECHO benchmark (Miglior et al., 2025), a task requiring long-range propagation. We compare EEGNN against APGCN (Spinelli et al., 2021), which also utilizes node-level exits. As shown in Table 4, EEGNN achieves state-of-the-art performance (MAE 0.065) by correctly maintaining deep exits. Figure 2 illustrates the exit statistics across 4 seeds: EEGNN consistently maintains exits in the deep regime, aligning more with the need for long-range reasoning. Conversely, APGCN exits prematurely due to its budget-aware bias, resulting in significantly higher error. Examples of task awareness in Appendix F.8.
Figure 2: Node exit level statistics on sssp: APGCN and EEGNN.
Task adaptivity. Second, we assess how EEGNN adapts depth to task definitions using Peptides-func (classification) and Peptides-struct (regression) from LRGB (Dwivedi et al., 2023). Despite sharing the same set of graphs, Figure 4 shows that EEGNN learns distinct exit distributions for each. For Peptides-func, the model predominantly exits after just two layers, while for Peptides-struct, it shifts to a slightly deeper distribution. This aligns with recent findings that these tasks do not strictly require long-range propagation (Bamberger et al., 2025). EEGNN’s adaptive shallow processing achieves $68\%$ AP, (see Table 18), demonstrating its ability to tailor computational depth to specific task requirements, a key advantage over fixed-depth GNNs. This experiments also complement those from ECHO, as here EEGNN learns to exit very early. Other robustness checks on the other datasets are provided in Appendix F.14.
Modularity. To illustrate modularity, we attach the exit module to standard MPNNs (Appendix F.6). We observe that in deep regimes ($L=20$), baseline MPNNs degrade severely, and early exits offer only unstable recovery. In contrast, EEGNN remains stable as depth grows, confirming that the SAS-GNN backbone provides the necessary stability to preserve information across layers, enabling robust early exiting (see Figures 1-3).
Runtime and efficiency. EEGNN reduces inference costs by skipping redundant computation, treating $L$ as a maximum budget rather than a fixed requirement. Table 2 (averaged over 1,500 passes, $m^{\prime}=32$) confirms that while baselines like Polynormer scale linearly with depth, EEGNN achieves nearly constant inference time even as $L$ increases. Furthermore, thanks to weight sharing, EEGNN maintains a constant parameter count significantly lower than competitive baselines, highlighting both computational and memory efficiency. Later in this section we discuss about the accuracy-efficiency trade-off of our models. Additional runtime and parameters analyses are provided in Appendix F.12.
Figure 3: Early-Exit on classic backbones presents unstable results on the Questions dataset.
@@
Figure 4: EEGNN’s Graph exit layer distributions. Numbers on top of the bars correspond to the number of test graphs that decided to exit at that layer. The associated performance is in Table 18.
Table 2: Runtime and parameter analysis in Questions. Times are in seconds.
@@
Model Inference Time (s) Number of Parameters 10 Layers 20 Layers 10 Layers 20 Layers GCN $0168\pm 0.0012 0.0$$0251\pm 0.0037$ 20,288 30,848 Co-GNN $0352\pm 0.0038 0.0$$0598\pm 0.0087$ 34,982 55,782 Polynormer $0278\pm 0.0053 0.0$$0381\pm 0.0078$ 50,404 80,100 SAS-GNN $0279\pm 0.0103 0.0$$0425\pm 0.0133$ 11,840 11,840 EEGNN $0216\pm 0.0064 0.0$$0257\pm 0.0078$ 12,562 12,562
@@
Table 3: Node classification under heterophily. The scores are marked in red for the first, blue for the second, and green for the third. Results are averaged across different graph splits. Scores are from Finkelshtein et al. (2024); Luo et al. (2024).